Let f(x) =


1 + 2x, x <= 0


3x - 2, 0 < x <= 1 2x ^ 2 - 1, x > 1


Check whether f is discontinuous. If yes, find where? ii) Give a rough sketch of the graph of f

1
Expert's answer
2021-11-02T15:02:29-0400

"f(x)=\begin{cases}
1+2x,\qquad x\leq 0
\\
3x-2,\qquad 0<x\leq 1
\\2x^2-1,\qquad x>1
\end{cases}"

1) "x=0"

"\lim\limits_{x\rightarrow 0-0}f( x)= \lim\limits_{x\rightarrow 0-0}(1+2x)=1"

"\lim\limits_{x\rightarrow 0+0}f( x)= \lim\limits_{x\rightarrow 0+0}(3x-2)=-2"


"\lim\limits_{x\rightarrow 0-0}f(x)\neq \lim\limits_{x\rightarrow 0+0}f(x)"


"f(x)" is discontinuous at "x=0"


2) "x=1"

"\lim\limits_{x\rightarrow 1-0}f( x)= \lim\limits_{x\rightarrow 1-0}(3x-2)=1"

"\lim\limits_{x\rightarrow 1+0}f( x)= \lim\limits_{x\rightarrow 1+0}(2x^2-1)=1"


"\lim\limits_{x\rightarrow 1-0}f(x)=\lim\limits_{x\rightarrow 1+0}f(x)=f(1)"


"f(x)" is continuous at "x=1"